A rhombus has sides of length 6 cm. One of its diagonals is 10 cm long. Find the length of the other diagonal ?

Answer by midwood_trail(310)

A rhombus has sides of length 6 cm. One of its diagonals is 10 cm long. Find the length of the other diagonal ?
The sides of a rhombus are the same measure.
One diagonal measures 10m. This means that it is 5m + 5m.
The two diagonals meet and form right angles.
We can use the Pythagorean Theorem to find one leg of one of the right triangles formed inside the rhombus.
5^2 + x^2 = 6^2
25 + x^2 = 36
x^2 = 36 - 25
x^2 = 11
Take the square root of both sides.
x = sqrt{11}.
The other diagonal will be sqrt{11} + sqrt{11) = 2sqrt{11}m
Did you follow?